OpenStudy (anonymous):

Translate each phrase and select the one that is DIFFERENT from the others. two more than a number, n the total of n and 2 twice the number n the sum of n and 2

OpenStudy (anonymous):

1) 2+n 2) x=n+2 3)2n 4)n+2=x

OpenStudy (anonymous):

thats what i take them as and the the one that would be different would be the 3rd one.